x/x+2 + 4/x-2 = 1
I need to solve this equation.

Answer by stanbon(75887)

(x/x+2) + (4/x-2) = 1
-----------------------------
Multiply thru by (x+2)(x-2)
x(x-2) + 4(x+2) = (x+2)(x-2)
x^2 -2x + 4x + 8 = x^2 - 4
2x = -12
x = -6
============================
Cheers,
Stan H.